Telephone company claims that the average length of local calls is 4.5 minutes.randome sample of 50 calls is observed and it was mean call length is 3.9 minutes with standard deviation of 2.8 minutes.Is sample results significantly different than claims if alpha is 0.010
1
Expert's answer
2020-08-24T08:08:23-0400

Given that the population mean "\mu" = 4.5 minutes as claimed by the company.


On observing a sample set of N = 50

sample mean is found to be "\bar x" = 3.9 and

standard deviation "\sigma" = 2.8


Let us assume that the null hypothesis H0 : "\mu" = 4.5 (company's claim is correct)


Alternate hypothesis will be Ha : "\mu \not = 4.5"


For the above given data we can calculate the value of Zcal = "\frac{\bar x - \mu}{\frac{\sigma}{\sqrt N}}"


Zcal = "\frac {3.9 - 4.5}{\frac{2.8}{\sqrt50}} = -1.515"


At the given level of significance "\alpha" = 0.01 we can find the value of Z("\frac {\alpha}{2}" ) from the standard normal table.


Since this is a two tailed test we consider "\alpha"/2 = 0.01/2 = 0.005


Therefore Z("\frac {\alpha}{2}" ) = Z0.005 = 2.5758


Since, |Zcal| = |-1.515| = 1.515 is found to be less than < Z("\frac {\alpha}{2}" ) = 2.5758,

we do not have sufficient evidence to reject the company's claim.


Hence we can say that the sample results are not significantly different than the company's claim.